I need help using the elimination method to solve the following:
2x+8y=12
-8x-8y=-16

Answer by stanbon(75887)

solve the following:
2x+8y=12
-8x-8y=-16
-------------
Divide thru 1st by 2
Divide thru 2nd by -8
----
x + 4y = 6
x + y = 2
----
Subtract and solve for "y":
3y = 4
y = 4/3
-----
Solve for "x":
x + y = 2
x + 4/3 = 2
x = 2/3
=====
Ans: (2/3 , 4/3)
====================
Cheers,
Stan H.
===============